OpenStudy (anonymous):

Identify the correct slope and y-intercept of the equation 2x + 4y = 8. @AriPotta I'm sorry Harry... can you just explain this to me? I know the Y and X axis already. y = (0,4) x = (2,0)

OpenStudy (anonymous):

lol Luis.

OpenStudy (aripotta):

rearrange the stuff to make it in slope-intercept form. so like, subtract 2x from both sides and then divide 4 from both sides

OpenStudy (aripotta):

then you can see the slope clearly

OpenStudy (aripotta):

or you could use the slope formula using those points

jimthompson5910 (jim_thompson5910):

First you need to solve for y \[\Large 2x + 4y = 8\] \[\Large 4y = -2x + 8\] \[\Large y = \frac{-2x + 8}{4}\] \[\Large y = \frac{-2x}{4} + \frac{8}{4}\] \[\Large y = -\frac{1}{2}x + 2\] The equation is now in y = mx + b form. This is slope-intercept form. In this case, m = -1/2 and b = 2 So the slope is -1/2 and the y-intercept as a point is (0, 2)

OpenStudy (anonymous):

Ohh, I see.... Thank you all C;

OpenStudy (anonymous):

'like' yeah. ;) Yes, I do get it... just a little rusty on slope. I did module 4 way back when and now I'm retaking the exam so.

OpenStudy (anonymous):

:)